# An Effective Version of the $p$-Curvature Conjecture for Order One Differential Equations

### v1

@@ 3.2 The Least Prime That Does Not Split in a Number Field
The matter of bounding the least rational prime number that does not split completely in a given (Galois) number field has been extensively studied in the literature. It is a special case of the more general problem of finding an upper bound on the least prime that has a prescribed splitting pattern in a number field [LMO79, AK19, KW22], in accordance with the Chebotarev Density Theorem. More precisely, given a Galois field extension $K/\mathbb{Q}$ of degree $d$ and discriminant $D$, there are several asymptotic estimates on the size of the smallest rational prime that does not split completely in $K$. An explicit statement was proven by Vaaler and Voloch [VV00, Thm. 1].
@@
Sharper, asymptotic bounds were proven, both under the assumption of the Generalized Riemann Hypothesis by Murty [Mur94], and unconditionally [Mur94, Li12].
@@
First, most results in this direction are of asymptotic nature, making it hard to convert them into an explicit bound. The exception is Vaaler and Voloch’s statement given above. However, they have an assumption on the minimal size of $D$ that exceeds the computational capacity of modern computers. Li claims [Li12, p. 1062] that this assumption is artificial and better results could be expected for small $D$.
@@ 4 An Effective Version of Kronecker’s Theorem
This section is devoted to the proof of Theorem 1.2. We outline the steps here. In [CC85] D.V. and G.V. Chudnovsky provide an elementary proof of Kronecker’s Theorem (and with it, according to the discussion in Section 2, of Honda’s proof of the $p$-curvature conjecture for order one equations) using Hermite-Padй approximation.
@@
The next ingredient is Hermite-Padй approximation. The ideas of this method were first introduced by Hermite in [Her73] to prove the transcendence of $e$, and then extended by him and Padй, his student [Her93, Pad92, Pad94]. It produces for a finite number of power series, a list of polynomials of prescribed degrees, such that the sum of the power series, weighted by the polynomials, has highest possible order. More precisely, let $K$ be a field. Let $f_{1}(z),\ldots,f_{m}(z)\in K[\![z]\!]$ be a tuple of power series. Let $n_{1},\ldots,n_{m}$ be non-negative integers and let $P_{1}(z),\ldots,P_{m}(z)\in K[z]$ be polynomials of degree $n_{1},\ldots,n_{m}$ respectively. If the power series
@@
has order greater than or equal to $n_{1}+\cdots+n_{m}+m-1$ at 0, then the tuple $(P_{1}(z),\ldots,P_{m}(z))$ is called a Hermite-Padй approximant to $(f_{1}(z),\ldots,f_{m}(z))$ of type $(n_{1},\ldots,n_{m})$, and $g(z)$ is called the remainder of the approximation. For further exposition of the concept, see [Mah68, Jag64, Jag64a].
In the following, a well-known formula for the Hermite-Padй approximants of the consecutive powers $(1-z)^{i\alpha}$ with $\alpha\not\in\mathbb{Q}$ for $1\leq i\leq 2M+1$, with coefficients of uniform degree $N$ for any integer values of $M$ and $N$ is exploited. It was introduced by Padй in [Pad01] and studied and extended by Jager, Mahler, and G. Chudnovsky [Jag64a, Mah68, Chu83], but could already be deduced from Hermite’s work [Her73]. The result also includes an explicit expression for the first non-vanishing term in the approximation.
@@
For the rest of the section we fix an irrational number $\alpha$ with annihilating polynomial $R(w)=r_{n}w^{n}+\cdots+r_{1}w+r_{0}$. We set $\Delta\coloneqq|r_{n}|$, and pick $B\in\mathbb{R}$ to be a bound on the maximal modulus of a root of $R(w)$. We let $L$ be the splitting field of $R$. The denominator $\operatorname{den}(\gamma)$ of $\gamma\in L$ is the smallest positive integer $a\in\mathbb{N}_{>0}$, such that $a\gamma\in\mathcal{O}_{L}$, the ring of integers of $L$. In particular, we have $\operatorname{den}(\alpha)|\Delta$. Let $M,N$ be two positive integers and define $\sigma=\sigma(M,N)\coloneqq(2M+1)N+2M$. We will assume that $R(w)$ splits into linear factors for all primes up to $\sigma$. The proof of Theorem 1.2 is equivalent to showing that when choosing $M\coloneqq\left\lceil 2.826\cdot r_{n}^{3}\cdot\delta(\Delta)\right\rceil$ and $N\coloneqq 6.076\cdot BM$ the above assumptions lead to a contradiction. We will accomplish this in Proposition 4.12, contradicting Corollary 4.10.
@@ 4.1 Hermite-Padй Approximants to Binomial Powers and Estimates on Their Remainder
Our definition of $\Omega$ differs slightly from the quantity called $\Omega$ in the Chudnovsky brothers’ article [CC85] because we believe that their line of argument, that we essentially followed and reproduced, does not hold with their choice. More precisely, the factors $(\pm k\alpha+N)$ do not appear in their definition of $\Omega$, hence for any $i\in\{1,\dots,2M+1\}$, the product $\Omega p_{i,N}$ simplifies to a polynomial in $\alpha$ with rational coefficients, times some factors of the form $(k\alpha+N)^{-1}$, for $k\in\mathbb{N}_{>0}$. The reasoning on the primes appearing in the denominator of this product does not hold anymore in this case (see the proof of Proposition 4.2, Equations (10) and (11)).
@@ 4.1.1 Bounding the Denominator
$\operatorname{den}(\Omega g_{\sigma})\leq C_{0}^{(2M-1)N}\Delta^{6MN+5M+N}\delta^{6MN+N+2M},$
where $C_{0}$ satisfies $\operatorname{lcm}(1,\dotsc,N)\leq C_{0}^{N}$ and $\delta=\prod\limits_{p|\Delta}p^{1/(p-1)}$.
@@
The existence of such a real number $C_{0}$ can be deduced from bounds on the Chebyshev function $\psi(x)=\log(\operatorname{lcm}(1,\dots,\lfloor x\rfloor))$. A possible $C_{0}$ is $\sqrt[113]{C}<2.826$, where $C$ is an explicit 51-digit number, see [RS62, Thm. 12] and OEIS A206431.
@@
Let $k,s\in\mathbb{Z}$, $r\in\mathbb{N}$, and let $p$ be a prime number. Let $\alpha$ be given as the zero of a polynomial $R(w)\in\mathbb{Z}[w]$ with leading coefficient $\Delta$.
@@
    If $p\not\hskip 2.5pt\mid\Delta$ and $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$, then $p\not\hskip 2.5pt\mid\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$.
@@
    If $p\not\hskip 2.5pt\mid\Delta$ and $p|\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$, then $p<r$ and $v_{p}\left(\binom{k\alpha+s}{r}\right)\geq-\left\lfloor\frac{r}{p-1}\right\rfloor$.
@@
    If $v_{p}(\Delta)=v$, then $v_{p}\left(\binom{k\alpha+s}{r}\right)\geq-\left(vr+\left\lfloor\frac{r}{p-1}\right\rfloor\right)$.
@@
    If $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$ for all primes $p\leq r$ not dividing $\Delta$, then $\operatorname{den}\left(\binom{k\alpha+s}{r}\right)$ divides $\Delta^{r}\prod_{p\mid\Delta}p^{\left\lfloor\frac{r}{p-1}\right\rfloor}.$
Доказательство.
Let $L/\mathbb{Q}$ be the splitting field of $R(w)$, and $\mathfrak{p}$ a prime ideal above $p$.
If $R(w)\bmod p$ splits in $\mathbb{F}_{p}[w]$, then there is a rational integer $a$ such that $\alpha\equiv a\bmod\mathfrak{p}$ hence $\binom{k\alpha+s}{r}\equiv\binom{ka+s}{r}\bmod\mathfrak{p}$ and the latter is a rational integer. Hence $v_{\mathfrak{p}}\left(\binom{k\alpha+s}{r}\right)\geq 0$. Denominators are rational integers and $\ell\in\mathbb{Z}$ belongs to $\mathfrak{p}$ if and only if $p|\ell$. Thus $p$ does not divide the denominator of $\binom{k\alpha+s}{r}$.
The second statement follows from the fact that $\binom{k\alpha+s}{r}=(k\alpha+s)_{r}/r!$ where $(u)_{r}\coloneqq u(u-1)\cdots(u-r+1)$ denotes the falling factorial. Thus the denominator of $\binom{k\alpha+s}{r}$ is composed of a product of primes dividing either $r!$ or the denominator of $\alpha$. Since $\operatorname{den}(\alpha)|\Delta$, we notice that $p\not\hskip 2.5pt\mid\operatorname{den}(\alpha)$ here. Primes dividing $r!$ are all primes smaller than $r$. Moreover it is a standard fact that $v_{p}(r!)\leq\frac{r}{p-1}$, hence the result.
For the third statement note that $p^{v}(k\alpha+s-\ell)$ has positive $p$-adic valuation for all integers $\ell$. Thus $v_{p}\left(p^{vr}(k\alpha+s)_{r}\right)\geq 0$ and we conclude as for the second statement.
@@
Remark 4.5.
  1. (1)
    The proof of Lemma 4.4, is independent of the choice of the prime ideal $\mathfrak{p}$ over $p$ since $L$ is Galois, as the splitting field of a polynomial.
  2. (2)
    In the setting of the second assertion, we know from the contrapositive of the first assertion that $R(w)\bmod p$ does not split completely in $\mathbb{F}_{p}[w]$.
@@
$\Delta^{4MN+3M}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{N}{p-1}\right\rfloor}\right)^{4M}\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
Moreover, $\operatorname{den}\left(\binom{(i-1)\alpha}{\sigma-h}\right)$ divides the expression $\Delta^{\sigma-h}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{\sigma-h}{p-1}\right\rfloor}\right),$ by Lemma 4.4 (4), which in turn divides $\Delta^{\sigma}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{\sigma}{p-1}\right\rfloor}\right).$
@@
$\Delta^{6MN+5M+N}\prod_{p|\Delta}\left(p^{\left\lfloor\frac{6MN+N+2M}{p-1}\right\rfloor}\right)\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
@@
$\operatorname{den}(\Omega g)\leq\Delta^{6MN+5M+N}\delta^{6MN+N+2M}C_{0}^{(2M-1)N},$
@@ 4.1.2 Bounding the Norm
$|\operatorname{Norm}(\Omega g_{\sigma})|^{1/d}<\frac{(2e^{2})^{M}\cdot\exp\left(\pi u_{1}(M)B+4\log\left(\frac{4}{3}\right)\frac{u_{2}(M)}{N}B^{2}+\frac{M+1}{6N}\right)}{(2M+1)^{(2M+1)N}\cdot N^{M}\cdot(2M+1)^{2M+1/2}\cdot\pi^{2M}},$
where $B\in\mathbb{N}$ is a bound on the maximal modulus of all roots of a given annihilating polynomial $R(w)$ of $\alpha$.
To obtain bounds on the factorials, we borrow the following variant of Stirling’s formula from [Rob55]: For $k\in\mathbb{N}_{>0}$ we have
@@ 4.2 A Bound on the Number of Primes
$\begin{array}[]{cl}X(M)&\coloneqq\cfrac{C_{0}^{2M-1}\Delta^{6M+1}\delta^{6M+1}}{(2M+1)^{2M+1}}=\cfrac{1}{C_{0}^{2}\Delta^{2}\delta^{2}}\cdot\left(\cfrac{C_{0}\Delta^{3}\delta^{3}}{2M+1}\right)^{2M+1}\\ Y(M,N)&\coloneqq\exp\left(\pi Bu_{1}(M)+4\log\left(\cfrac{4}{3}\right)B^{2}\cfrac{u_{2}(M)}{N}\right)\cdot(2M+1)^{-2M-1/2}\cdot\left(\cfrac{2e^{2}\Delta^{5}\delta^{2}}{\pi^{2}}\right)^{M}\\ Z(M,N)&\coloneqq\left(\cfrac{1}{N}\right)^{M}\cdot\exp\left(\cfrac{M+1}{6N}\right).\end{array}$
Of course, these quantities also depend on $\Delta$, which is suppressed in the notation. Moreover, once we fix $M$ to be a function of $\Delta$, the notation becomes particularly misleading. However, in the following calculations it will be convenient to view these quantities as functions of $M$ and $N$.
@@
From the asymptotics of these bounds, one can easily deduce the existence of parameters $M$ and $N$, such that $|\operatorname{den}(\Omega g_{\sigma})\operatorname{Norm}_{L/\mathbb{Q}}(\Omega g_{\sigma})^{\frac{1}{d}}|<1$. For large enough $M$ compared to $\Delta$, the expression $X(M)$ is strictly smaller than $1$. Moreover, if we choose $N$ large enough with respect to $M$, we have $X(M)^{N}<Y(M,N)^{-1}$, as the right-hand side is independent of $N$. Finally, for $N>M>3$, clearly $Z(M,N)<1$. Thus, for this choice of $M,N$ we find the desired contradiction to the assumption that $\alpha$ was irrational, and so we have proved Kronecker’s Theorem. This argumentation was carried out by the Chudnovsky brothers in [CC85]. The rest of this subsection is devoted to showing that our explicit choices of $M$ and $N$ suffice for the contradiction to occur.

### v2

@@ 3.2 The Least Prime That Does Not Split in a Number Field
The matter of bounding the least rational prime number that does not split completely in a given (Galois) number field has been extensively studied in the literature. It is a special case of the more general problem of finding an upper bound on the least prime that has a prescribed splitting pattern in a number field [26, 1, 22], in accordance with the Chebotarev Density Theorem. More precisely, given a Galois field extension $K/\mathbb{Q}$ of degree $d$ and discriminant $D$, there are several asymptotic estimates on the size of the smallest rational prime that does not split completely in $K$. An explicit statement was proven by Vaaler and Voloch [50, Thm. 1].
@@
Sharper, asymptotic bounds were proven, both under the assumption of the Generalized Riemann Hypothesis by Murty [32], and unconditionally [32, 27].
@@
First, most results in this direction are of asymptotic nature, making it hard to convert them into an explicit bound. The exception is Vaaler and Voloch’s statement given above. However, they have an assumption on the minimal size of $D$ that exceeds the computational capacity of modern computers. Li claims [27, p. 1062] that this assumption is artificial and better results could be expected for small $D$.
@@ 4 An Effective Version of Kronecker’s Theorem
This section is devoted to the proof of Theorem 1.2. We outline the steps here. In [11] D.V. and G.V. Chudnovsky provide an elementary proof of Kronecker’s Theorem (and with it, according to the discussion in Section 2, of Honda’s proof of the $p$-curvature conjecture for order one equations) using Hermite-Padй approximation.
@@
The next ingredient is Hermite-Padй approximation. The ideas of this method were first introduced by Hermite in [16] to prove the transcendence of $e$, and then extended by him and Padй, his student [17, 34, 35]. It produces for a finite number of power series, a list of polynomials of prescribed degrees, such that the sum of the power series, weighted by the polynomials, has highest possible order. More precisely, let $K$ be a field. Let $f_{1}(z),\ldots,f_{m}(z)\in K[\![z]\!]$ be a tuple of power series. Let $n_{1},\ldots,n_{m}$ be non-negative integers and let $P_{1}(z),\ldots,P_{m}(z)\in K[z]$ be polynomials of degree $n_{1},\ldots,n_{m}$ respectively. If the power series
@@
has order greater than or equal to $n_{1}+\cdots+n_{m}+m-1$ at 0, then the tuple $(P_{1}(z),\ldots,P_{m}(z))$ is called a Hermite-Padй approximant to $(f_{1}(z),\ldots,f_{m}(z))$ of type $(n_{1},\ldots,n_{m})$, and $g(z)$ is called the remainder of the approximation. For further exposition of the concept, see [29, 20, 21].
In the following, a well-known formula for the Hermite-Padй approximants of the consecutive powers $(1-z)^{i\alpha}$ with $\alpha\not\in\mathbb{Q}$ for $1\leq i\leq 2M+1$, with coefficients of uniform degree $N$ for any integer values of $M$ and $N$ is exploited. It was introduced by Padй in [36] and studied and extended by Jager, Mahler, and G. Chudnovsky [21, 29, 12], but could already be deduced from Hermite’s work [16]. The result also includes an explicit expression for the first non-vanishing term in the approximation.
@@
For the rest of the section we fix an irrational number $\alpha$ with square-free annihilating polynomial $R(w)=r_{n}w^{n}+\cdots+r_{1}w+r_{0}$ with $r_{n}>0$. We set $\Delta_{R}\coloneq\operatorname{res}_{w}(R(w),R^{\prime}(w))\neq 0$, we define $\delta(\Delta_{R})\coloneqq\prod_{p|\Delta_{R}}p^{1/(p-1)},$ and we pick $B\in\mathbb{R}$ to be a bound on the maximal modulus of a root of $R(w)$. We let $L$ be the splitting field of $R$. The denominator $\operatorname{den}(\gamma)$ of $\gamma\in L$ is the smallest positive integer $a\in\mathbb{N}_{>0}$, such that $a\gamma\in\mathcal{O}_{L}$, the ring of integers of $L$. In particular, we have $\operatorname{den}(\alpha)|r_{n}$. Let $M,N$ be two positive integers and define $\sigma=\sigma(M,N)\coloneqq(2M+1)N+2M$. We will assume that $R(w)$ splits into linear factors for all primes up to $\sigma$, that do not divide $r_{n}$. The proof of Theorem 1.2 is equivalent to showing that when choosing $M\coloneqq\left\lceil 2.826\cdot r_{n}^{3}\cdot\delta(\Delta_{R})\right\rceil$ and $N\coloneqq 6.076\cdot BM$ the above assumptions lead to a contradiction. We will accomplish this in Proposition 4.12, contradicting Corollary 4.10.
@@ 4.1 Hermite-Padй Approximants to Binomial Powers and Estimates on Their Remainder
Our definition of $\Omega$ differs slightly from the quantity called $\Omega$ in the Chudnovsky brothers’ article [11] because we believe that their line of argument, that we essentially followed and reproduced, does not hold with their choice. More precisely, the factors $(\pm k\alpha+N)$ do not appear in their definition of $\Omega$, hence for any $i\in\{1,\dots,2M+1\}$, the product $\Omega p_{i,N}$ simplifies to a polynomial in $\alpha$ with rational coefficients, times some factors of the form $(k\alpha+N)^{-1}$, for $k\in\mathbb{N}_{>0}$. The reasoning on the primes appearing in the denominator of this product does not hold anymore in this case (see the proof of Proposition 4.2, Equations (10) and (11)).
@@ 4.1.1 Bounding the Denominator
$\operatorname{den}(\Omega g_{\sigma})\leq C_{0}^{(2M-1)N}r_{n}^{6MN+5M+N}\delta^{6MN+N+2M},$
where $C_{0}$ satisfies $\operatorname{lcm}(1,\dotsc,N)\leq C_{0}^{N}$ and $\delta=\prod\limits_{p|\Delta_{R}}p^{1/(p-1)}$.
@@
The existence of such a real number $C_{0}$ can be deduced from bounds on the Chebyshev function $\psi(x)=\log(\operatorname{lcm}(1,\dots,\lfloor x\rfloor))$. A possible $C_{0}$ is $\sqrt[113]{C}<2.826$, where $C$ is an explicit 51-digit number, see [41, Thm. 12] and OEIS A206431.
@@
Let $k,s\in\mathbb{Z}$, $t\in\mathbb{N}$, and let $p$ be a prime number. Let $\alpha$ be given as the zero of a square-free polynomial $R(w)\in\mathbb{Z}[w]$ with leading coefficient $r_{n}$ and $\Delta_{R}\coloneqq\operatorname{res}_{w}(R(w),R^{\prime}(w))$.
@@
    If $p\not\hskip 2.5pt\mid\Delta_{R}$ and $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$, then $p\not\hskip 2.5pt\mid\operatorname{den}\left(\binom{k\alpha+s}{t}\right)$.
@@
    If $p\not\hskip 2.5pt\mid r_{n}$ and $p|\operatorname{den}\left(\binom{k\alpha+s}{t}\right)$, then $p\leq t$ and $v_{p}\left(\binom{k\alpha+s}{t}\right)\geq-\left\lfloor\frac{t}{p-1}\right\rfloor$.
@@
    If $v_{p}(r_{n})=v$, then $v_{p}\left(\binom{k\alpha+s}{t}\right)\geq-\left(vt+\left\lfloor\frac{t}{p-1}\right\rfloor\right)$.
@@
    If $R(w)\bmod p$ splits completely in $\mathbb{F}_{p}[w]$ for all primes $p\leq t$ not dividing $r_{n}$, then $\operatorname{den}\left(\binom{k\alpha+s}{t}\right)$ divides $r_{n}^{t}\prod_{p\mid\Delta_{R}}p^{\left\lfloor\frac{t}{p-1}\right\rfloor}.$
Proof.
If $R(w)\bmod p$ splits in $\mathbb{F}_{p}[w]$, and $p\not\hskip 2.5pt\mid\Delta_{R}$ then all of its zeroes are distinct elements of $\mathbb{F}_{p}$. Then, by Hensel’s Lemma [33, (4.6) p.129] $R(w)$ splits over $\mathbb{Z}_{p}$ into linear factors, i.e. $\alpha\in\mathbb{Z}_{p}$. Thus for any $j\geq 1$ there exists an integer $0\leq a_{j}<p^{j}$, such that $\alpha\equiv a\bmod p^{j}$. Hence, $\binom{k\alpha+s}{t}=\binom{ka_{j}+s}{t}+b_{j}\frac{p^{j}}{t!}$ for some $p$-adic integer $b_{j}$. Thus for large enough $j$, we obtain $(ka_{j}+s)_{t}\neq 0$, where $(u)_{t}\coloneqq u(u-1)\cdots(u-t+1)$ denotes the falling factorial, and $v_{p}\left(\binom{k\alpha+s}{t}\right)\geq v_{p}\left(\binom{ka_{j}+s}{t}\right)\geq 0$.
The second statement follows from the fact that $\binom{k\alpha+s}{t}=(k\alpha+s)_{t}/t!$. Thus the denominator of $\binom{k\alpha+s}{t}$ is composed of a product of primes dividing either $t!$ or the denominator of $\alpha$. Since $\operatorname{den}(\alpha)|r_{n}$, we notice that $p\not\hskip 2.5pt\mid\operatorname{den}(\alpha)$ here. Primes dividing $t!$ are all primes smaller than $t$. Moreover it is a standard fact that $v_{p}(t!)\leq\frac{t}{p-1}$, hence the result.
For the third statement note that $p^{v}(k\alpha+s-\ell)$ has positive $p$-adic valuation for all integers $\ell$. Thus $v_{p}\left(p^{vt}(k\alpha+s)_{t}\right)\geq 0$ and we conclude as for the second statement.
@@
Example 4.5.
We illustrate that the condition $p\not\hskip 2.5pt\mid\Delta_{R}$ in (1) that ensures the existence of a Hensel lift is necessary. Let $R(w)\coloneqq w^{2}-2\cdot w+2$ with leading coefficient $r_{2}=1$, and $\Delta_{R}=-4$. Let $\alpha=i+1$ be its root. Pick $p=2\mid\Delta_{R}$. Then $\binom{i+1}{2}=\frac{i-1}{2}$ is not $2$-integral.
@@
$r_{n}^{4MN+3M}\prod_{p|\Delta_{R}}\left(p^{\left\lfloor\frac{N}{p-1}\right\rfloor}\right)^{4M}\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
Moreover, $\operatorname{den}\left(\binom{(i-1)\alpha}{\sigma-h}\right)$ divides the expression $r_{n}^{\sigma-h}\prod_{p|\Delta_{R}}\left(p^{\left\lfloor\frac{\sigma-h}{p-1}\right\rfloor}\right),$ by Lemma 4.4 (4), which in turn divides $r_{n}^{\sigma}\prod_{p|\Delta_{R}}\left(p^{\left\lfloor\frac{\sigma}{p-1}\right\rfloor}\right).$
@@
$r_{n}^{6MN+5M+N}\prod_{p|\Delta_{R}}\left(p^{\left\lfloor\frac{6MN+N+2M}{p-1}\right\rfloor}\right)\left(\operatorname{lcm}_{h\in\{0,\ldots,N\}}\binom{N}{h}\right)^{2M-1}.$
@@
$\operatorname{den}(\Omega g)\leq r_{n}^{6MN+5M+N}\delta^{6MN+N+2M}C_{0}^{(2M-1)N},$
@@ 4.1.2 Bounding the Norm
$|\operatorname{Norm}(\Omega g_{\sigma})|^{1/d}<\frac{(2e^{2})^{M}\cdot\exp\left(\pi u_{1}(M)B+4\log\left(\frac{4}{3}\right)\frac{u_{2}(M)}{N}B^{2}+\frac{M+1}{6N}\right)}{(2M+1)^{(2M+1)N}\cdot N^{M}\cdot(2M+1)^{2M+1/2}\cdot\pi^{2M}}.$
To obtain bounds on the factorials, we borrow the following variant of Stirling’s formula from [40]: For $k\in\mathbb{N}_{>0}$ we have
@@ 4.2 A Bound on the Number of Primes
$\begin{array}[]{cl}X(M)&\coloneqq\cfrac{C_{0}^{2M-1}r_{n}^{6M+1}\delta^{6M+1}}{(2M+1)^{2M+1}}=\cfrac{1}{C_{0}^{2}r_{n}^{2}\delta^{2}}\cdot\left(\cfrac{C_{0}r_{n}^{3}\delta^{3}}{2M+1}\right)^{2M+1}\\ Y(M,N)&\coloneqq\exp\left(\pi Bu_{1}(M)+4\log\left(\cfrac{4}{3}\right)B^{2}\cfrac{u_{2}(M)}{N}\right)\cdot(2M+1)^{-2M-1/2}\cdot\left(\cfrac{2e^{2}r_{n}^{5}\delta^{2}}{\pi^{2}}\right)^{M}\\ Z(M,N)&\coloneqq\left(\cfrac{1}{N}\right)^{M}\cdot\exp\left(\cfrac{M+1}{6N}\right).\end{array}$
Of course, these quantities also depend on $r_{n}$, and $\Delta_{R}$, which are suppressed in the notation. Moreover, once we fix $M$ to be a function of $r_{n},\Delta_{R}$, the notation becomes particularly misleading. However, in the following calculations it will be convenient to view these quantities as functions of $M$ and $N$.
@@
From the asymptotics of these bounds, one can easily deduce the existence of parameters $M$ and $N$, such that $|\operatorname{den}(\Omega g_{\sigma})\operatorname{Norm}_{L/\mathbb{Q}}(\Omega g_{\sigma})^{\frac{1}{d}}|<1$. For large enough $M$ compared to $r_{n},\Delta_{R}$, the expression $X(M)$ is strictly smaller than $1$. Moreover, if we choose $N$ large enough with respect to $M$, we have $X(M)^{N}<Y(M,N)^{-1}$, as the right-hand side is independent of $N$. Finally, for $N>M>3$, clearly $Z(M,N)<1$. Thus, for this choice of $M,N$ we find the desired contradiction to the assumption that $\alpha$ was irrational, and so we have proved Kronecker’s Theorem. This argumentation was carried out by the Chudnovsky brothers in [11]. The rest of this subsection is devoted to showing that our explicit choices of $M$ and $N$ suffice for the contradiction to occur.